OpenStudy (joftheworld):

I need help finding three points to this linear equation please.. -x+2y=2

OpenStudy (crashonce):

substitute any three numbers just do x=1, x=2, x=3 and you will get three numbers for y. they are your three points CO

OpenStudy (joftheworld):

can you give an example on the first part ?

OpenStudy (crashonce):

when x=1: -1+2y=2 therefore 2y=3 therefore y=1.5 therefore one point is (1,1.5)

OpenStudy (joftheworld):

that's another problem .. im not allowed to use decimals ..

OpenStudy (crashonce):

then you can just use 3/2

OpenStudy (joftheworld):

or fractions.. sorry for all the work.. it just confuses me

OpenStudy (crashonce):

1.5=3/2 im not sure how else you can express it

OpenStudy (joftheworld):

I have to graph it, and im not allowed to put the decimal in between numbers..

OpenStudy (crashonce):

in these questions, we generally use fractions in graphs because they are more accurate then decimals

OpenStudy (crashonce):

oh to make it easier, just do x=2, x=4, x=6 then you can plot on the grid easier

OpenStudy (joftheworld):

okay ill try

OpenStudy (joftheworld):

wait so then it'll be -(2)+2y=2 ? and so on? (4) (6)

OpenStudy (crashonce):

yes

OpenStudy (joftheworld):

so would it be (-6,8) (-4,6) (-3,6) ?

OpenStudy (joftheworld):

woops last one should be a -2

OpenStudy (crashonce):

i got: (2,1), (4,3), (4,6)

OpenStudy (crashonce):

nop (2,2 should bbe first

OpenStudy (joftheworld):

oh okay thanks ! do you think you can help me with a few more ?

OpenStudy (crashonce):

sure

OpenStudy (joftheworld):

-x-y=1

OpenStudy (crashonce):

subs any three values ill get (0,-1), (1,-2), (2, -3)

OpenStudy (joftheworld):

okay thanks, and x+y=-2 ?

OpenStudy (crashonce):

(0,-2), (1, -3), (2-4)

OpenStudy (joftheworld):

thank you so much, also -3y=-2x-6

OpenStudy (crashonce):

(0,2), (3, 4), (6, 6) tricky one, im trying to get whole numbers no problems

OpenStudy (crashonce):

in th substituting in numbers to get your values

OpenStudy (joftheworld):

okay thanks can you help me with 5 more questions ?

OpenStudy (crashonce):

just post them in one go

OpenStudy (joftheworld):

2y=3x+11

OpenStudy (crashonce):

(0, 5.5) (1,7), (2,8.5) ok i gtg post them and ill get back to you later

OpenStudy (joftheworld):

okay thanks for the help !